–3(10x – 3) + 10(3x – 3) = x – 7
Answer by rapaljer(4551)

–3(10x – 3) + 10(3x – 3) = x – 7
-30x + 9 + 30x - 30 = x-7
-21= x-7

Add + 7 to each side:
-21+7= x - 7 + 7
-14 = x

R^2 at SCC